# Algebraic Connectivity Enhances Hyperedge Specificity in the Alzheimer’s Disease Continuum

### v1

@@ 4 Discussion
The hypergraph is a data structure able to model high-order relations across brain regions, thus better approximating and representing a more realistic nature of the brain. The key concepts of the hypergraph are related to the hyperedges and their weights, since they represent groups of highly related brain regions and the importance of the different hyperedges in the hypergraph, respectively.
@@
Selecting the most appropriate method for computing the hyperedge weight is essential, since it carries the functional information that resides in the brain areas of that subgraph. Assigning constant values or setting all weights to ones can be too simplistic yang2021graphlshc ; bai2021hypergraph and should be avoided since no meaningful features can be extracted. Similarly, assigning the hyperedge degree as the weight pisarchik2025hypergraph results in a partial loss of the regionals’ functional data, keeping only the information related to the number of regions forming the hyperedges.
@@
In this study, taking advantage of the brain functional data, more complex approaches have been considered and developed for computing the hyperedges’ weights. Specifically, we explored approaches such as the Gaussian similarity kernel niu2023applications ; peng2023multiview ; gao2020hypergraph ; zhang2017joint , the L2 norm of LASSO coefficients balogh2022hypergraph ; gao2020hypergraph ; pan2021characterization , the mean or sum of pairwise correlations values between time series of the regions forming each hyperedge bai2025hypergraph ; song2023brain ; shao2023analysis , and the mean correlation values of the regions defined by the Schaefer functional atlas. With these approaches, we identified the hyperedges 9 (somatomotor areas) and 75 (somatomotor and ventral attention regions) and three FNs (somatomotor, salience, ventral attention, and default mode) that showed significant differences in AD versus MCI and AD versus HC, highlighting hipoconnectivity in the patients’ groups with respect to the HC one. These findings were consistent across the baseline approaches, including the outcomes of the Schaefer FNs, and with previous works where hipoconnectivity was observed in AD wiesman2021somatosensory ; li2012attention . On the other hand, an hyperconnectivity effect was featured in the hyperedge 59 (somatomotor and ventral attention areas) found by the L2 norm-based method, where higher connectivity was found in AD.
@@
Even if these approaches obtained relevant results concordant with previous works, they relied on averaging or summing pair-wise metrics, thus not featuring the real nature of the hypergraph, except for the L2 norm applied to the LASSO coefficients that aggregated the regression coefficients. On these bases, the advantage of using the $a(\mathcal{G})$ derives from the fact that featuring the Laplacian matrix of the subgraph (hyperedge) and its decomposition to the second-smallest eigenvalue, we can extract a measure of how well the hyperedges are connected, and avoiding the use of mean of sum operations which brings approximation in the final results. The proposed method identified a larger number of hyperedges (9, 22, 48, 60, 75, and 77, where 9 and 75 were in common with the other approaches) statistically different across groups. These hyperedges included regions of the somatomotor, dorsal and ventral attentional, default mode, and visual networks. Interestingly, some hyperedges (e.g., 48, 60, and 75) grouped regions belonging to different FNs, such as somatomotor and attentional, and visual and default mode, which are usually considered separately in standard atlases. This suggests that our approach with hypergraphs could capture more insights and functional organization in AD that conventional methods overlook. Most of the significant hyperedges showed reduced $a(\mathcal{G})$ in AD patients, consistent with the idea of a disease-related loss of functional integration pini2021breakdown . In contrast, hyperedges involving dorsal and ventral attentional networks showed increased connectivity in the AD patients, reflecting possible compensatory effects. Further, the functional information that resided in regions belonging to the somatomotor and ventral attention FNs (hyperedges 60 and 75) showed a partial mediation between entorhinal tau load and cognitive performance, indicating that tau pathology may impair cognition by disrupting the coordination of large-scale networks in AD iordan2024salience ; sperling2019impact ; bejanin2017tau ; mielke2017association ; jack2016t . A similar pattern was previously observed employing the FC networks related to episodic memory and executive function as mediators between the tau levels and the memory and executive scores. They found that a significant mediation effect of FC was present between tau and memory function, but not in the executive ziontz2024behaviorally . Their results partially agree with our findings where for the hyperedges 60 and 75, significant mediation effects were present for memory score. On the contrary, only in the hyperedge 75 was a mediation effect between tau and executive function observed due to FC. Taken together, the obtained results highlighted the plausibility and potential of the proposed approach for deriving hyperedge weights that can capture meaningful functional information in the AD continuum and help understanding the neurophysiological mechanisms underlying Alzheimer’s Disease.
@@ 4.1 Main Limitations
One possible limitation of this work could be related to the LASSO method employed for computing the hypergraph structure. Even if this method has been widely used and the objective of this work was not to define new strategies for extracting the hypergraph structure, more performant approaches than LASSO could be considered for defining the hypergraph structure. Another possible limitation could be related to the $a(\mathcal{G})$ computation. This approach requires the graph to be connected, otherwise, it will result in a null score, thus limiting this approach only to connected graphs.
@@ 5 Conclusion
In this work, we proposed an approach for computing the hyperedge weights in the hypergraph structure relying on the algebraic connectivity $a(\mathcal{G})$, which is a measure of how well the brain regions composing the hyperedges are connected, even if they reside in different FNs. The results achieved with the $a(\mathcal{G})$ method showed a greater number of significantly different hyperedges than common approaches. The combination of decreased connectivity in sensorimotor, attentional, and default mode networks, together with increased connectivity in attention systems, underscores the heterogeneous and dynamic nature of the disease. Our use of $a(\mathcal{G})$ and hypergraph models provides a new way to capture these changes, highlighting both functional disconnection and potential adaptations. Furthermore, the mediation effects involving tau burden reinforce the idea of FC as the mechanistic link between molecular pathology and clinical symptoms, offering a pathway through which disease processes affect cognition.
@@ Acknowledgments
This study was funded by the Ministero dell’Università e della Ricerca (Bando PRIN 2022, “AI4BRAVE: AI for modeling of the Brain-Heart Axis in aging” project-reference code 202292PHR2), by the Ministero dell’Istruzione e del Merito (MIUR D.M. 737/2021, “AI4Health: empowering neurosciences with eXplainable AI methods” project), and by FSE project ANTIAGING: “Artificial intelligence for healthy aging” (project-reference code B31J23000910002) as well as NIH grant # R01AG073949 and # 1R01AG090597, and NSF grant # 2112455.
Additional information
The data used in this work were collected by ADNI (https://adni.loni.usc.edu/), and they are publicly available after requested on the ADNI website.

### v2

@@ 4 Discussion
The hypergraph is a data structure able to model high-order relations across brain regions, thereby better approximating and representing a more realistic nature of the brain. The key concepts of the hypergraph are related to the hyperedges and their weights, since they represent groups of highly related brain regions and the importance of the different hyperedges in the hypergraph, respectively.
@@
Selecting the most appropriate method for computing the hyperedge weight is essential, since it carries the functional information that resides in the brain areas of that subgraph. Assigning constant values or setting all weights to ones can be too simplistic Yang and others (2021); Bai et al. (2021) and should be avoided since no meaningful features can be extracted. Similarly, assigning the hyperedge degree as the weight Pisarchik et al. (2025) results in a partial loss of the regional functional data, keeping only the information related to the number of regions forming the hyperedges.
@@
In this study, taking advantage of the brain functional data, more complex approaches have been considered and developed for computing the hyperedges’ weights. Specifically, we explored approaches such as the Gaussian similarity kernel Niu and Du (2023); Peng et al. (2023); Gao et al. (2020); Zhang et al. (2017), the L2 norm of LASSO coefficients Balogh et al. (2022); Gao et al. (2020); Pan et al. (2021), the mean or sum of pairwise correlations values between time series of the regions forming each hyperedge Bai et al. (2025); Song et al. (2023); Shao et al. (2023), and the mean correlation values of the regions defined by the Schaefer 100 functional atlas. With these approaches, we identified the hyperedges 9 (somatomotor areas) and 75 (somatomotor and salience/ventral attention regions) and three FNs (somatomotor, salience, salience/ventral attention, and default-mode) that showed significant differences in AD vs MCI and AD vs HC, highlighting hypoconnectivity in the patients’ groups with respect to the HC one. These findings were consistent across the baseline approaches, including the outcomes of the Schaefer 100 FNs, and with previous works that have observed hypoconnectivity in AD Wiesman and others (2021); Li et al. (2012). On the other hand, a hyperconnectivity effect was featured in the hyperedge 59 (somatomotor and salience/ventral attention areas) found by the L2 norm-based method, where higher connectivity was found in AD.
@@
Even if these approaches obtained relevant results concordant with previous works, they relied on averaging or summing pairwise metrics, thus not featuring the real nature of the hypergraph, except for the L2 norm applied to the LASSO coefficients that aggregated the regression coefficients. On these bases, the advantage of using the $a(\mathcal{G})$ derives from the fact that featuring the Laplacian matrix of the subgraph (hyperedge) and its decomposition to the second-smallest eigenvalue, we can extract a measure of how well the hyperedges are connected avoiding the use of mean or sum operations which brings approximation in the final results. The proposed method identified a larger number of hyperedges (9, 22, 48, 60, 75, and 77, where 9 and 75 were in common with the other approaches) statistically different across groups. These hyperedges included regions of the somatomotor, dorsal and salience/ventral attention, default-mode, and visual FNs. Interestingly, some hyperedges (e.g., 48, 60, and 75) grouped regions belonging to different FNs, such as somatomotor and attention, and visual and default-mode, which are usually considered separately in standard atlases. This suggests that our approach with hypergraphs could capture more insights and functional organization in AD that conventional methods overlook. Most of the significant hyperedges showed reduced $a(\mathcal{G})$ in AD patients, consistent with the idea of a disease-related loss of functional integration Pini et al. (2021). This pattern appeared most prominently in hyperedges formed by regions belonging to the salience/ventral attention, somatomotor, visual, and default-mode FNs which have been previously linked to AD in earlier studies Dolci et al. (2025a, b); Dai et al. (2019). Specifically, prior works have observed functional disruptions in the salience/ventral attention and somatomotor FNs, indicating alterations in these systems during AD progression. Likewise, reduced functional integrity in the default-mode and visual FNs has been associated with AD, with greater declines in these FNs corresponding to more severe cognitive impairment Dai et al. (2019); Huang et al. (2021); Agosta et al. (2010); Li et al. (2015); Jalilianhasanpour et al. (2019). In contrast, hyperedges involving dorsal and salience/ventral attention FNs showed increased connectivity in the AD patients, reflecting possible compensatory effects. The compensatory effect pattern in these FNs is still debated in the literature. However, some studies found a similar trend to the one observed in our results, i.e., AD greater than MCI greater than HC in terms of functional power, thus confirming our proposed results Franzmeier et al. (2017); Jalilianhasanpour et al. (2019); Balthazar et al. (2014).
Further, $a(\mathcal{G})$ demonstrated a higher capability to address various classification tasks using the extracted features, i.e., the hyperedge weights. $a(\mathcal{G})$ outperformed the other approaches in the three classification tasks except for the most critical one (i.e., HC vs MCI), where similar performances to the Gaussian similarity kernel were obtained. Regarding the importance of the extracted features, hyperedge 75 was the most relevant hyperedge that guided the classifications, consistent with the larger effect size found across groups, in contrast to the other hyperedge weights.
Additionally, the relationship between tau pathology and cognitive impairment was partially mediated by the high-order functional information resided in hyperedges 60 and 75, suggesting possible modulations in the coordination of large-scale networks in ADIordan and others (2024); Sperling and others (2019); Bejanin and others (2017); Mielke and others (2017); Jack Jr and others (2016). A similar pattern was previously observed employing the FC networks related to episodic memory and executive function as mediators between the tau levels and the memory and executive scores. They found that a significant mediation effect of FC was present between tau and memory function, but not in the executive Ziontz and others (2024). Their results partially agree with our findings, where, for the hyperedges 60 and 75, significant mediation effects were present for the memory score. On the contrary, only in the hyperedge 75 a mediation effect between tau and executive function due to FC was observed.
The potential of $a(\mathcal{G})$ was further supported using a higher-resolution atlas and an independent cohort. In general, $a(\mathcal{G})$ demonstrated stronger discriminative power in pairwise comparisons and across groups compared to the other methods, except for the Gaussian similarity kernel. Indeed, the Gaussian similarity kernel detected one additional statistically different hyperedge when an atlas at a higher resolution was employed. The biological composition of the hyperedges found by $a(\mathcal{G})$, considering both Schaefer 400 and an independent ADNI cohort, revealed FNs (salience/ventral and dorsal attention, default-mode, somatomotor, and visual) that were consistent with the results of Schaefer 100. Considering even higher resolution atlases or additional datasets with different acquisition parameters (e.g., scan length and TR) could be of interest for assessing further the stability of the framework and thus deserve further investigations, even if some attentions have to be taken. In particular, fine-grained atlases (e.g., 800-1000 parcels) could hamper a direct biological interpretation of the information that resides in the small parcels, and might introduce noisy features with possible unstable hyperedges.
Taken together, the obtained results highlighted the plausibility and potential of the $a(\mathcal{G})$ for deriving hyperedge weights that can capture meaningful functional information in the AD continuum and help understand the neurophysiological mechanisms underlying AD.
@@ 4.1 Main Limitations
One possible limitation of this work could be related to the LASSO method employed for computing the hypergraph structure. Even if this method has been widely used and the objective of this work was not to define new strategies for extracting the hypergraph structure, other approaches than LASSO could be considered for defining it. Another possible limitation could be related to the $a(\mathcal{G})$ computation. This approach requires the graph to be connected, otherwise it will result in a null score. Such a constraint limits this approach only to connected graphs. Even though no disconnected subgraphs were obtained in the current study, a feasible strategy is to discard those hyperedges from the analysis since no comparisons could be performed. Finally, even if $a(\mathcal{G})$ showed better results than the other methods, the Cliff’s $\delta$ effect size found across the three groups was relatively small.
@@ 5 Conclusion
In this work, we proposed an approach for computing the hyperedge weights in the hypergraph structure relying on the $a(\mathcal{G})$, which is a measure of how well the brain regions composing the hyperedges are connected, even if they reside in different FNs. The results obtained with $a(\mathcal{G})$ showed a greater number of significantly different hyperedges and superior classification performance than common approaches. The combination of decreased connectivity in somatomotor, attention, and default-mode FNs, together with increased connectivity in attention systems, underscores the heterogeneous and dynamic nature of the disease. Our use of $a(\mathcal{G})$ and hypergraph models provides a new way to capture these changes, highlighting both functional disconnection and potential adaptations. Furthermore, the mediation effects involving tau burden reinforce the idea of FC as the mechanistic link between molecular pathology and clinical symptoms, offering a pathway through which disease processes affect cognition.
Data Availability Statement
The data used in this work were collected by ADNI (https://adni.loni.usc.edu/), and they are publicly available upon request on the ADNI website.
The code used for the analyses of this work can be found in this GitHub repository: https://github.com/BraiNavLab/Algebraic_Connectivity_Hypergraph_Alzheimer.git.